OpenStudy (adunb8):

differential & linear algebra question !!! 3AM stress!

OpenStudy (adunb8):

\[2t(d^2x/dt^2) = (dx/dt)^2 \]

OpenStudy (adunb8):

have to turn it into 1st order diff equation.

OpenStudy (anonymous):

ISn't it quite evident - you have derivative-of-a-derivative. Sooo if u want 1-st order choose as your variable\[u = \frac{ dx }{ dt }\]

OpenStudy (anonymous):

??

OpenStudy (adunb8):

i see i just got that part and got up to \[\int\limits_{}^{} 1/(v^2-1) dv = \int\limits_{}^{}(1/2t) dt\]

OpenStudy (anonymous):

seems Too final - can you show the intermediate stages ?

OpenStudy (anonymous):

It is unclear to me where did you get the "1" in the left hand integral

OpenStudy (anonymous):

@adunb8 ?

OpenStudy (adunb8):

okay hold

OpenStudy (adunb8):

i forgot that there was -1 at the right side

OpenStudy (anonymous):

So my guidance was useful. what do you say ?

OpenStudy (adunb8):

yea but the problem is im not getting my teacher's answer.. its suppose to be \[x=-t-2/cln[ct-1]+k\]

OpenStudy (anonymous):

You did half the soln. you found x' - now integrate that once more. And - if you are normal human, put a medal where medal is due.

OpenStudy (adunb8):

im stuck here ... not getting the answer...

OpenStudy (adunb8):

i did integrated and got ln |v^2-1| = ln|2t| then cancelled each out using e^ and left with v^2-1 = 2t + C

OpenStudy (anonymous):

on the left hand side you have to do partial fraction like this\[\frac{ 1 }{ \left( u-1 \right)\left( u+1 \right) }=\frac{ A }{ \left( u-1 \right) }+\frac{ B }{ \left( u+1 \right) }\]then solving for A and B and continue do integral you'll get two terms of ln

OpenStudy (anonymous):

It is all small technical steps - do them with precision.

OpenStudy (adunb8):

oh i see partial fraction ??

OpenStudy (anonymous):

the final solution in the term of u, you'll get this\[\frac{ u+1 }{ u-1}=ct\]then de integral angain, since you've defined u = dx/dt

OpenStudy (adunb8):

dx/dt + 1 / dx/dt -1

OpenStudy (anonymous):

\[\int\limits_{}^{}dx=-\int\limits_{}^{}\frac{ \left( 1+ct \right) }{ \left( 1-ct \right) }dt\]

OpenStudy (anonymous):

It's OK for you now ??

OpenStudy (adunb8):

yeah thank for the clear instruction =)

OpenStudy (anonymous):

no problem sir :)

OpenStudy (anonymous):

I like you @Pinguine

OpenStudy (adunb8):

me to i love you both =)

OpenStudy (anonymous):

><"

OpenStudy (anonymous):

much ♥♥♥♥♥♥♥♥♥♥♥♥♥♥♥♥♥ love ♥♥♥♥♥♥♥♥♥everybody♥♥♥♥♥♥♥♥